OpenStudy (anonymous):

(3)/(5)/(1)/(5). I think this is divide? Should be fractions

OpenStudy (anonymous):

3/5 / 1/5 = 3/5 * 5/1 =15/5 =3

OpenStudy (anonymous):

Its simple when we have to divide a fraction by a fraction, we just multiply the numerator by the reciprocol of the denominator.

OpenStudy (anonymous):

OK. thanks, that looks easy enough.

OpenStudy (anonymous):

It needs to be pointed out that the expression (3)/(5)/(1)/(5) is highly troublesome. It is supposed to be interpreted as (((3 divided by 5) divided by 1) divided by 5) based on the typical order of operations. There is a significant difference between (3)/(5)/(1)/(5) and ((3)/(5))/((1)/(5)). So, what exactly did you mean? Shubham assumed that it was ((3)/(5))/((1)/(5)). Is he correct in this presumption? If so, it looks like \[ \frac{3/5}{1/5}? \]